a 40,000 kg airplane is provided with 130,000N of thrust by its engines as it takes of at an angle of 25 degrees above the horizontal. If the velocity of the airplane is constant, how much upward force on it is provided by its vertical component


1
Expert's answer
2020-11-23T10:31:35-0500

Explanations & Calculations


  • Thrust imposed on the rocket is along the moving direction.
  • Vertical component of the thrust is given by "\small130000\sin\theta=\small130000\sin25=\small54940.37N".
  • Therefore, the upward force is 54940.37N


  • There is a mistake that 40000 should be the weight of the rocket not mass "\small 40000kg\to40000N" because if it is mass, the thrust won't lift the rocket.
  • Since the rocket is moving at a uniform velocity, there are no any net forces acting on it, rather forces are in equilibrium.
  • The difference in the vertical components is equal to the resistive forces acting on the rocket.

"\small {54940.37-40000=14940.37N}"